OpenStudy (anonymous):

carol is twice lorraine's age. in six years, the sum of their ages will be 30. how old is each now?

OpenStudy (anonymous):

Lorraine's age = L Carol's age now = 2L <---"twice Carol's age. Lorrains's age in 6 years = (L + 6) Carol's are in 6 years = (2L + 6) <--- ie Carol's age now plus 6 years. Put both equations to get the sum of their ages in 6 years: \[(2L+6)+(L+6) = 30\] Solve for L to get Lorraine's age right now. Derive Carol's age from Lorraine's age now.